# PACE: Pose Annotations in Cluttered Environments

### v1

@@ 4 Dataset Statistics
In pursuit of diversity, we placed the objects ten disparate environments, each featuring varying levels and configurations of occlusion. We document 10 video sequences per scene, each sequence encompassing 1 to 5 objects. We captured RGB and depth data using an Intel RealSense D415 camera, with a resolution of . The capture process spans distances ranging from 0.5m to 1.5m from the objects.
4.1 Object Distribution and Diversity
The comprehensive distribution of objects is depicted in Figure 3, showcasing a diverse array of both rigid and articulated models. As demonstrated in Figure 5, the dataset encompasses a broad spectrum of object sizes, with the majority measuring approximately 0.2m along the bounding box diagonal. This includes larger items such as storage bins, adding to the heterogeneity of the dataset.
4.2 Variability in Pose, Occlusion, and Environmental Context
The dataset is characterized by a rich variability in object poses, as illustrated in Figure 6, which outlines the statistical distribution of azimuth and elevation angles, indicating comprehensive spatial coverage.
Moreover, we categorized and analyzed the occlusion levels, as shown in Figure 7, delineating instances of severe, moderate, and minor occlusions. Such classification is crucial for assessing the robustness of pose estimation algorithms against varying degrees of visibility.
A visual representation of the ten distinct environmental settings used for data capture is provided in Figure 8, reflecting the contextual diversity of the dataset.
4.3 Dataset Split for Training, Validation, and Testing
To facilitate a comprehensive and equitable evaluation of pose estimation methodologies, we partition the dataset into validation and test subsets following a 20/80 ratio. This separation ensures that a substantial volume of data is available for rigorous testing.
Additionally, to support research necessitating extensive training datasets, we generate a considerable set of synthetic images utilizing a physically based renderer (PBR) [6]. The training set comprises over 52,000 images, each featuring multiple objects, culminating in a vast number of annotations. Qualitative examples of the synthetic images are presented in Figure 9.
@@ 5 Evaluation Benchmarks
State-of-the-art pose estimation algorithms typically decompose the task into two distinct stages: detecting or segmenting the object of interest, followed by pose estimation within the predicted bounding box or mask. While the latter has been the primary focus in literature, often leveraging off-the-shelf detectors, a comprehensive evaluation requires an isolated assessment of each stage. Therefore, we examine the pose estimation result under the assumption of perfect detection to allow for equitable comparison across methods. The object detection performance of current state-of-the-art methods can be found in the supplementary and is not the focus of this dataset. We also introduce a benchmark for object pose tracking, premised on the availability of ground-truth data in the initial video frame.
@@ 5.1 Pose Estimation Benchmark
This benchmark is divided into instance-level and category-level pose estimation. The former concerns known instances during training, while the latter involves unknown instances similar to those in the training set.
@@ 5.1.1 Instance-Level Pose Estimation
Adhering to the BOP challenge protocol [15], we utilize Average Recall (AR) of Visible Surface Discrepancy (VSD), Maximum Symmetry-Aware Surface Distance (MSSD), and Maximum Symmetry-Aware Projection Distance (MSPD) as our metrics. Detailed computation of these metrics is described in [15]. Objects with visibility fraction less than are skipped for evaluation following the BOP [15] convention. Results are averaged across all instances.
@@
We assess four baselines: PPF [8], CosyPose [19], SurfEmb [12], and GDRNPP [40]. While PPF does not require training data, the others are state-of-the-art methods dependent on additional training on the PBR dataset. To focus on pose estimation performance, we assume that ground-truth instance detections are available.
@@
The quantitative analysis is presented in Table 3. Notably, while state-of-the-art methods excel on the BOP benchmarks, they do not perform as well as PPF, which relies on local geometric feature matching in point clouds. The relative success of PPF highlights the potential for improvement in state-of-the-art methods, particularly regarding robustness and scalability to handle a large set of instances. Qualitative results can be found in the supplementary.
| Modality | Detection | AR | AR | AR | AR | |
| PPF [8] | D | G.T. | 35.3 | 42.7 | 49.3 | 42.4 |
@@
| Detection | IoU | IoU | AP | ||||||
| 0:20 | 0:60 | 0:5cm | 0:15cm | 0:20 0:5cm | 0:60 0:15cm | ||||
| NOCS [41] | Mask-RCNN | 0.0/0.0 | 0.0/0.0 | 0.0/0.0 | 0.1/0.0 | 2.8/0.0 | 26.1/0.0 | 0.0/0.0 | 0.1/0.0 |
| HS-Pose [47] | G.T. | 32.7/0.2 | 7.3/0.0 | 5.2/0.0 | 7.0/0.4 | 61.4/41.0 | 86.2/79.5 | 4.0/0.0 | 6.6/0.4 |
| SGPA [4] | G.T. | 1.3/0.0 | 0.0/0.0 | 6.7/1.1 | 13.4/9.0 | 18.5/10.4 | 58.4/49.2 | 3.3/0.3 | 11.7/7.7 |
| DualPoseNet [25] | G.T. | 0.1/0.0 | 0.0/0.0 | 5.1/0.0 | 5.3/0.0 | 15.0/34.6 | 53.6/69.4 | 1.4/0.0 | 4.1/0.0 |
| SAR-Net [24] | G.T. | 25.3/0.5 | 1.1/0.0 | 5.2/0.1 | 7.1/2.6 | 37.6/37.1 | 77.3/77.9 | 2.8/0.1 | 6.3/2.6 |
@@ 5.1.2 Category-Level Pose Estimation
We adopt from NOCS [41], calculating mean Average Precision (AP) across predefined angle and translation thresholds. Specifically, AP@0:20 represents AP averaged from 0 to 20 at 1 intervals; AP@0:5cm is averaged from 0cm to 5cm at 0.25cm intervals; AP@0:20,0:5cm combines these angle and translation thresholds. IoU and IoU denote AP for 3D bounding box matches at IoU thresholds of 25 and 50, respectively. Objects with visibility fraction less than are skipped for evaluation following the BOP [15] convention. Results are separately reported for both rigid and articulated objects.
@@
We evaluate six recent category-level pose estimation methods: NOCS [41], HS-Pose [47], SGPA [4], DualPoseNet [25], SAR-Net [24], and ANCSH [22], the latter specifically designed for articulated objects. For adapting rigid-object-focused methods to articulated objects, each movable part is considered a distinct category. For instance, scissors with two movable parts are treated as two separate categories. Ground-truth detections are presumed for each method, except for NOCS, which outputs both instance masks and pose estimations using a unified network. The compared methods all use the depth as input.
@@
Quantitative comparisons are listed in Table 4. On rigid objects, HS-Pose shows excellent performance on IoU and IoU, as well as AP metrics for translation, but falls short on rotation AP metrics compared to SGPA, suggesting a proficiency in size and translation prediction but not rotation. SGPA excels in rotation estimation but demonstrates lower performance on bounding box size metrics. For articulated objects, all methods face challenges due to the movable parts, with ANCSH underperforming on the large-scale real-world dataset, indicative of a significant sim-to-real gap. NOCS shows instability, likely due to its reliance on RGB-based coordinate prediction, which lacks robustness and scalability. Qualitative results are available in supplementary.
@@ 5.2 Pose Tracking Benchmark
We categorize state-of-the-art pose tracking methods into model-free, which use a 3D CAD model, and model-based, requiring only the initial pose at the first frame.
@@
For model-free pose tracking, we use four metrics following previous work [39]. 1) , the percentage of predictions with rotation error 5° and translation error 5cm. 2) IoU25, the percentage of intersection over union that is larger than between the two 3D bounding boxes with ground-truth size, transformed by the predicted and ground-truth 6D pose, respectively. 3) , mean value of rotation error in degrees. 4) , mean value of translation error in centimeters. Here the last two metrics are respect to IoU25 since objects with IoU are not counted.
For model-based pose tracking, we report the area under curve (AUC) with respect to ADD, ADD-S [45] and ADD(-S). We set the maximum threshold of AUC to be 0.1m [38]. The ADD metric is first introduced in [14] to calculate average per-point distance between two point clouds, transformed by the predicted pose and the ground-truth, respectively. For symmetric objects like bowls, ADD-S metric is introduced to count for the point correspondence ambiguity. The notation ADD(-S) corresponds to computing ADD for non-symmetric objects and ADD-S for symmetric objects. The objects with visibility fraction less than are skipped for evaluation following the BOP [15] convention. Results are reported by averaging over the rigid/articulated categories, separately.
@@
| Training-Free | Modality | 55cm | IoU25 | R | T | |
@@
For methods relying on object mask during tracking, we directly used the ground-truth mask. Since the big gap of performance between rigid and articulated objects, we separately reported their average results.
We regarded each part of the articulated objects as independent for all tracking methods except for CAPTRA [44] which treat all parts of an object as a whole. We trained CAPTRA for each category using our synthetic PBR data.
Result Analysis:
Tables 5 and 6 present the performance metrics of state-of-the-art (SOTA) tracking methods on the PACE dataset with both rigid and articulated results separated by slash. The methods exhibit limited success, with ICG [34] achieving the highest AUC with respect to ADD(-S) at only 38.1% for rigid objects and a mere 10.1% for articulated objects, as shown in Table 5. Table 6 underscores the challenge further, with the best 55cm accuracy below 13%, and IoU25 not exceeding 47%, indicating substantial pose estimation errors for over half of the objects.
The convention followed by 6-PACK [39] and BundleTrack [42] disregards objects with IoU when computing rotational and translational errors (R and T). However, in the context of our dataset where methods generally struggle, this approach could mask true performance levels. This discrepancy is exemplified by BundleTrack [42], which reports lower performance in 55cm and IoU25 metrics yet shows seemingly better results for R and T. Such results indicate that current SOTA methods may require significant improvements to handle the complexity presented by the PACE dataset effectively. Qualitative results can be found in the supplementary.
6 Conclusions and Future Work
This work presented a comprehensive benchmark for 3D object pose estimation and tracking through the introduction of the PACE dataset. Our findings reveal that while there have been significant advancements in pose estimation techniques, there exists a substantial performance gap when these methods are applied to real-world, diverse datasets such as PACE. Particularly, current state-of-the-art methods struggle with articulated objects and exhibit limitations in robustness and scalability. This benchmark serves not only as a testament to the progress achieved but also as a clarion call for the research community to address the complexities of real-world applications.
The results from the PACE dataset underscore a pronounced generalization gap, suggesting that existing models may not sufficiently capture the complexities inherent in diverse real-world scenarios. To bridge this gap, future research could explore the potential of larger, more complex models for pose estimation that can encapsulate a wider variety of object features and environmental contexts.
@@ 7 Acknowledgements
This work was supported by the National Key Research and Development Project of China (No. 2021ZD0110700), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), Shanghai Qi Zhi Institute, and SHEITC (2018-RGZN-02046). Yang You is also supported in part by the Outstanding Doctoral Graduates Development Scholarship of Shanghai Jiao Tong University.
Supplementary Material
A Object Detection Benchmark
Though the primary focus of this dataset is the pose estimation, we are interested in how current state-of-the-art detection models perform on our dataset. In practice, the best method can be served as the default detection method when evaluating the pose estimation result.
A.1 Instance-Level Object Detection
Evaluation Metrics:
Consistent with established literature, we employ Average Precision (AP), Average Precision at IoU thresholds of 50% (AP) and 75% (AP), along with Average Recall (AR) as the metrics for evaluation.
Baseline Methods:
The task is categorized into bounding box (BBox) detection and instance segmentation (Mask). We benchmark the performance against three models: Mask R-CNN [13], YOLO-X [10], and MaskDINO [20]. MaskDINO is notable for its claim of providing precise bounding box and instance segmentation results. Models are trained on a provided PBR dataset, treating each instance as an individual ”category“, culminating in 576 unique categories. Results are averaged across all instances in the test set.
Results and Analysis:
Quantitative outcomes are presented in Table 7. We see that YOLO-X is the best in terms of the bounding box detection while MaskDINO is the best in instance segmentation.
A.2 Category-Level Object Detection
Evaluation Metrics:
The metrics for category-level object detection are identical to those used for instance-level detection.
Baseline Methods:
This task is bifurcated into bounding box (BBox) detection and instance segmentation (Mask). We evaluate the performance of YOLO-X [10], MaskDINO [20], and the zero-shot detector GLIP [21]. GLIP is capable of inferring bounding boxes based on textual descriptions of target categories. For instance segmentation, we compare Mask R-CNN [13], MaskDINO [20], and Grounded SAM [28], which can generate instance masks from text prompts in a zero-shot fashion. Non-zero-shot methods are trained on images from the provided PBR training set.
Results and Analysis:
Quantitative findings are detailed in Table 8. YOLO-X again is the winner is detecting bounding boxes and MaskDINO is the winner in instance segmentation. Though zero-shot methods can achieve some results but they are still far behind the supervised methods, partially due to the text ambiguity in describing the objects.
| Type | ZS | AP | AP | AP | AR | |
| YOLO-X [10] | BBox | ✗ | 50.9 | 66.4 | 57.9 | 64.3 |
| MaskDINO (BBox) | BBox | ✗ | 46.8 | 65.3 | 53.3 | 59.7 |
| GLIP [21] | BBox | ✓ | 22.2 | 30.3 | 26.1 | 59.8 |
| Mask R-CNN [13] | Mask | ✗ | 40.9 | 60.3 | 47.8 | 51.4 |
| MaskDINO (Mask) | Mask | ✗ | 42.9 | 65.1 | 47.4 | 54.6 |
| Grounded-SAM [28] | Mask | ✓ | 8.2 | 11.7 | 9.6 | 15.0 |
B Baseline Adaptations
This section delineates the adaptations made to the baseline methodologies enabling their evaluation on our dataset. Unless otherwise stated, the configurations adhere to the defaults specified in their respective originating papers.
B.1 Handling Object Symmetry
For baselines tasked with rotational regression, we address the ambiguity presented by object symmetry, where multiple rotations correspond to a single input by normalizing these rotations into distinct, unambiguous targets, following the method outlined in [31].
B.2 Instance-Level Pose Estimation Baselines
CosyPose [19]
To maintain comparability, we exclude the random background pasting augmentation from CosyPose. Moreover, to accommodate our hardware constraints, we reduce the image resolution by half.
SurfEmb [12]
With SurfEmb, images are decoded into a shared continuous embedding space alongside 3D point embeddings. Due to the extensive variety of objects in our dataset (576 in total), a separate image decoder for each would exceed our GPU’s memory capacity. Consequently, we implement a unified decoder across all objects.
B.3 Pose Tracking Baselines
CAPTRA [44]
We treat all objects as non-symmetric for training purposes, given that CAPTRA’s loss is designed to handle only asymmetry or continuous symmetry along the y-axis. We also limit our scope to articulated objects comprising two parts, aligning with CAPTRA’s fixed-part count assumption. To manage the higher image resolution and variable object size in our dataset compared to NOCS, we decrease the radius parameter when computing the axis-aligned bounding box aabb and downsample the back-projected point clouds prior to ball filtering.
C Qualitative Comparison on Pose Estimation/Tracking
In this section, we present additional illustrations that further elucidate the performance of various baseline methodologies. Specifically, Figure 10 provides a visualization of the efficacy of instance-level pose estimation methods. Similarly, Figure 11 demonstrates the performance of category-level pose estimation techniques. Additionally, Figure 12 and Figure 13 offer visual representations of the performance metrics for model-based and model-free tracking methods, respectively. These visualizations serve to complement the quantitative analyses provided earlier, offering a more comprehensive understanding of each method’s effectiveness.
D Annotation Software and Pipeline Details
Our annotation interface features a primary display that simultaneously presents views from a triple-camera setup. Adjacently, a sidebar is integrated, showcasing the current annotations within the scene, along with a catalog of potential 3D models from our database. Users can initiate the pose estimation by aligning corresponding points on the 3D models with their 2D image counterparts, utilizing a RANSAC-PnP algorithm. Subsequent refinements to the pose are facilitated through incremental rotations and translations, which can be executed via keys.
For enhanced user experience and efficiency, we have designed a tripartite panel system. This allows users to seamlessly toggle between 2D annotation, 3D annotation, and 2D segmentation workflows. As illustrated in Figure 14, this configuration enables real-time previews of annotations in both 2D and 3D, ensuring an intuitive and dynamic annotation process.
E Snapshots for All the Objects
Figure 15 gives the model snapshots for all our collected objects.

### v3

@@ 4 Dataset Statistics
In pursuit of diversity, we placed the objects in ten disparate environments, each featuring varying levels and configurations of occlusion. We captured 10 video sequences per scene, each sequence including 1 to 7 objects. We captured RGB and depth data using an Intel RealSense D415 camera, with a resolution of . The capture process spans distances ranging from 0.5m to 1.5m.
4.1 Annotation Distribution and Object Diversity
The comprehensive distribution of pose annotation counts is depicted in Figure 5, showcasing a diverse array including both rigid and articulated objects. As demonstrated in Figure 7, our object database encompasses a broad spectrum of object sizes, with the majority measuring near 0.2m along the bounding box diagonal. It also includes larger items such as storage bins, adding to the heterogeneity of the dataset.
4.2 Variability in Occlusion, Pose, and Environmental Context
The dataset has rich variability in occlusion levels and object poses, as illustrated in Figure 7, which outlines the statistical distribution of both visibility ratios and azimuth/elevation angles, indicating a comprehensive coverage reflecting real-world scenarios. These statistics can be helpful for assessing the robustness of pose estimation algorithms. Besides, the ten distinct capturing environments are shown in Figure 8, illustrating diverse backgrounds and lighting conditions.
4.3 Simulated Training Data and Dataset Split
To facilitate a comprehensive and equitable evaluation of pose estimation methodologies, we partitioned our entire real-world dataset into validation and test subsets following a 20/80 ratio. The test subset is used for both instance-level and category-level pose estimation evaluation tasks. The real-world validation subset is used for instance-level pose estimation only. We render another photo-realistic validation subset for category-level pose estimation, in order to cover enough category-level objects that are different from the test set.
Additionally, to support research necessitating extensive training datasets, we generated a large synthetic dataset called PACE-Sim that contains two different training subsets, for instance-level and category-level pose estimation, respectively. We use a physically based renderer [6] to ensure the images are highly photo-realistic. The synthetic training sets consist of 100K images, each featuring dozens of objects, yielding 2.4M annotations. The distribution of object instances in both PACE and PACE-Sim can be found in supplementary.
@@ 5 Evaluation Benchmarks
We assess both pose estimation and tracking. State-of-the-art (SOTA) pose estimation techniques generally involve two phases: first detecting or segmenting the object, and then estimating its pose within the identified mask. For a fair comparison, we evaluate pose estimation assuming perfect object detection. Detailed SOTA detection performance is available in the supplementary materials but is beyond this dataset’s primary scope.
For pose tracking, we provide ground-truth data in the initial frame and require methods to accurately track the object’s rotations and translations through the video sequence. Both model-free and model-based trackers are assessed.
@@ 5.1 Pose Estimation Benchmark
This benchmark is divided into instance-level and category-level pose estimation. The former concerns object instances seen during training (in different scenes), while the latter concerns instances which were not seen but come from the same category. We will also give an in-depth analysis of SOTA’s failure modes.
@@ 5.1.1 Instance-Level Pose Estimation
Adhering to the BOP challenge protocol [14], we use Average Recall (AR) of Visible Surface Discrepancy (VSD), Maximum Symmetry-Aware Surface Distance (MSSD), and Maximum Symmetry-Aware Projection Distance (MSPD) as our metrics. Detailed computation of these metrics is described in [14]. Results are averaged across all instances.
@@
We assessed four baselines: PPF [8], CosyPose [19], SurfEmb [12], and GDRNPP [33]. While PPF does not require training data, the others are SOTA methods dependent on additional training on the PBR dataset. To focus on pose estimation performance, we assume that ground-truth detections are available.
@@
The quantitative analysis is presented in Table 2. Notably, while SOTA methods excel on the BOP benchmarks, they do not perform as well as PPF, which relies on local geometric feature matching in point clouds. The relative success of PPF highlights the potential for improvement in SOTA methods, particularly regarding sim-to-real ability and scalability to handle a large set of instances. Qualitative results can be found in the supplementary.
| Modality | Detection | ARVSD | ARMSSD | ARMSPD | AR | |
| PPF [8] | D | G.T. | 53.4 | 48.1 | 55.2 | 52.2 |
@@
| Detection | IoU25 | IoU50 | AP | ||||||
| 0:20∘ | 0:60∘ | 0:5cm | 0:15cm | 0:20∘ 0:5cm | 0:60∘ 0:15cm | ||||
| NOCS [34] | Mask-RCNN | 0.2/0.0 | 0.0/0.0 | 1.2/0.0 | 4.3/0.0 | 17.0/0.0 | 33.5/0.0 | 1.1/0.0 | 4.2/0.0 |
| HS-Pose [43] | G.T. | 36.6/0.0 | 2.7/0.0 | 5.3/0.0 | 8.6/0.3 | 48.7/34.1 | 81.0/75.6 | 3.9/0.0 | 8.1/0.3 |
| SGPA [4] | G.T. | 2.6/0.0 | 1.2/0.0 | 5.6/1.1 | 11.4/7.3 | 16.1/6.8 | 50.7/17.5 | 3.3/0.9 | 10.1/7.2 |
| DualPoseNet [22] | G.T. | 24.2/0.3 | 0.1/0.0 | 5.5/0.0 | 8.1/0.0 | 18.6/27.3 | 63.4/65.2 | 1.8/0.0 | 6.5/0.0 |
| SAR-Net [21] | G.T. | 22.8/0.1 | 0.2/0.0 | 5.3/0.0 | 8.8/0.8 | 48.6/38.5 | 79.8/69.7 | 3.7/0.0 | 8.2/0.8 |
| CPPF++ [39] | G.T. | 44.5/3.6 | 4.4/0.0 | 15.2/1.7 | 27.3/6.2 | 35.3/31.1 | 74.0/66.7 | 9.9/1.1 | 24.9/5.9 |
| ANCSH [20] | G.T. | -/0.0 | -/0.0 | -/0.0 | -/0.2 | -/18.6 | -/50.4 | -/0.0 | -/0.2 |
@@ 5.1.2 Category-Level Pose Estimation
We adopted from NOCS [34], calculating mean Average Precision (AP) across predefined angle and translation thresholds. Specifically, AP@0:20∘ represents AP averaged from 0∘ to 20∘ at 1∘ intervals; AP@0:5cm is averaged from 0cm to 5cm at 0.25cm intervals; AP@0:20∘,0:5cm combines these angle and translation thresholds. IoU25 and IoU50 denote AP for 3D bounding box matches at IoU thresholds of 25 and 50, respectively. Results are separately reported for both rigid and articulated objects.
@@
We evaluated seven recent category-level pose estimation methods: NOCS [34], HS-Pose [43], SGPA [4], DualPoseNet [22], SAR-Net [21], CPPF++ [39] and ANCSH [20], the latter specifically designed for articulated objects. For adapting rigid-object-focused methods to articulated objects, each movable part is considered a distinct category. Ground-truth detections are presumed for each method, except for NOCS, which outputs both instance masks and pose estimations using a unified network.
@@
Quantitative comparisons are listed in Table 3. On rigid objects, HS-Pose and SAR-Net show fair performance on AP metrics for translation, but falls short on rotation AP metrics compared to CPPF++, suggesting a proficiency in size and translation prediction but not rotation. CPPF++ shows the best performance but is weaker on translation estimation. For articulated objects, all methods face challenges due to the movable parts, with ANCSH underperforming on the large-scale real-world dataset, indicative of a significant sim-to-real gap. NOCS shows instability, likely due to its inaccurate mask prediction. Qualitative results are available in supplementary.
5.1.3 Failure Mode Analysis
This section delves into the analysis of two critical questions that underline the failure of SOTA deep learning-based methods in pose estimation: 1) Is the failure attributed to the sim-to-real gap, considering the exclusive use of synthetic RGB-D images for training? 2) Does the failure stem from the inherent large-scale complexity of the PACE dataset, which poses a significant challenge for baseline models?
Ablation Setup:
To explore these questions, we introduced a new 80/10/10 train/validation/test split for the real-world data, diverging from the 0/20/80 split described in Section 4.3. We then examined the following settings:
Real2Real Single Instance/Category Fitting: This basic setup involves training on a single instance or category selected (we randomly select the category can and sample a can with object ID 57) from the 80/10/10 split of real data.
Sim2Real Single Instance/Category Fitting: This scenario differs from the first by utilizing all simulated data for training and validation, with evaluation on the same real-world test split.
Real2Real All Instance/Category Fitting: Unlike the first setting, this involves training, validation, and testing across all instances or categories.
| Training Setting | Instance-Level (AR%) | Category-Level (mean AP%@0:60∘, 0:15cm) | |||||||
| CosyPose | SurfEmb | GDRNPP | NOCS | HS-Pose | SGPA | DualPoseNet | SAR-Net | CPPF++ | |
| Real2Real Single Inst./Cat. | 76.4 | 86.7 | 88.2 | 76.9 | 81.2 | 86.7 | 58.6 | 59.3 | 83.4 |
| Sim2Real Single Inst./Cat. | 67.1 (-9.3) | 73.7 (-13.0) | 78.1 (-10.1) | 55.8 (-21.1) | 1.1 (-80.1) | 0.3 (-86.4) | 1.4 (-57.2) | 33.2 (-26.1) | 55.8 (-27.6) |
| Real2Real All Inst./Cat. | 8.2 (-68.2) | 11.2 (-75.5) | 9.7 (-78.5) | 0.3 (-76.6) | 5.6 (-75.6) | 9.8 (-76.9) | 1.2 (-57.4) | 2.1 (-57.2) | 13.7 (-69.7) |
Result Analysis and Discussions:
Our analysis in Table 4 reveals that baseline models effectively handle single instance or category fitting under the Real2Real Single Instance/Category Fitting setting, provided the training and test distributions align. However, the sim-to-real gap, especially noticeable in depth-input methods like DualPoseNet, HS-Pose, and SGPA, highlights significant challenges, as these methods struggle with the depth noise in real world.
Additionally, learning across all instances/categories leads to a notable performance drop, as shown in Table 4. This drop, together with suboptimal training data performance and loss plateauing, suggests underfitting and highlights the complexity posed by the PACE dataset’s diversity. Innovative approaches or larger foundational models might be necessary to address these challenges and effectively bridge the sim-to-real gap.
@@ 5.2 Pose Tracking Benchmark
We distinguish SOTA pose tracking techniques into model-based, utilizing a 3D CAD model, and model-free, which only require an initial pose from first frame.
@@
For model-based tracking, we measure the area under curve (AUC) for ADD, ADD-S [38], and ADD(-S) metrics, with a maximum threshold of 0.1m [32]. The ADD metric, introduced in [13], computes the average per-point distance between point clouds transformed by the predicted and ground-truth poses. ADD-S accounts for point correspondence ambiguities in symmetric objects.
For model-free tracking, we adopt four metrics from prior studies [31]: 1) , indicating the percentage of predictions within 5° rotation and 5cm translation errors. 2) IoU25, the percentage of instances where the intersection over union of the 3D bounding boxes exceeds . 3) and 4) , the mean rotational and translational errors in degrees and centimeters, respectively, only considering instances with IoU25%. We exclude objects with visibility under 10% as per the BOP convention [14]. Results are averaged separately across rigid and articulated object categories.
@@
| Training-Free | Modality | 5∘5cm | IoU25 | Rerr | Terr | |
@@
For tracking methods using object masks, we applied the ground-truth masks directly. Except for CAPTRA, which also predicts the 3D bounding box size, all methods estimate only rotation and translation; however, we used ground-truth bounding box sizes for consistency in evaluations. Unlike other methods that treat each part of an articulated object independently, CAPTRA considers the entire object as one entity, trained on our synthetic PBR data.
Result and Failure Mode Analysis:
Tables 5 and 6 display the performance of state-of-the-art (SOTA) tracking methods on the PACE dataset, highlighting both rigid and articulated object results. The highest achieved ADD(-S) for model-based methods is only 38.1% for rigid and 10.1% for articulated objects by ICG. Furthermore, Table 6 reveals severe limitations of model-free methods, with the best 5∘5cm accuracy under 13%, posing major tracking challenges.
The convention followed by 6-PACK [31] and BundleTrack [35] disregards objects with IoU when computing rotational and translational errors. However, this approach could overestimate the performance on Rerr and Terr. BundleTrack [35] reports lower performance in 5∘5cm and IoU25 metrics yet shows seemingly better results for Rerr and Terr, suggesting that it can only track well within a threshold but cannot recover once the object is lost. Qualitative results can be found in the supplementary.
6 Conclusions and Limitations
We present PACE, a new large-scale benchmark for 3D object pose estimation and tracking, designed with various occlusion levels across ten diverse environments. This benchmark both reflects prior progress and urges further research to tackle real-world complexities. Our analysis indicates significant advances in pose estimation, yet underscores a notable real-world performance gap, particularly with articulated objects and in terms of robustness and scalability. Future efforts could focus on advanced sim-to-real techniques and developing larger models that capture a broader spectrum of object features and environments.
Limitations
Due to resource limitations, PACE lacks large and valuable objects like tables, cameras, and laptops, which may affect its breadth in representing real-world object scenarios.
@@ 7 Acknowledgements
This work is supported by the National Key Research and Development Project of China (No. 2022ZD0160102), National Key Research and Development Project of China (No. 2021ZD0110704), Shanghai Artificial Intelligence Laboratory, XPLORER PRIZE grants. Yang You is also supported in part by the Outstanding Doctoral Graduates Development Scholarship of Shanghai Jiao Tong University.